# Near-linear Time Dispersion of Mobile Agents

### v2

@@ 4.1 Overview
A level-up event requires one leader $a$ and one zombie $b$ with the same level. That zombie $b$ will get level 0. Thereafter, $b$ never triggers a level-up event again because the level of a leader is monotonically non-decreasing starting from level 1. Therefore, for any $i\geq 1$, the number of agents that can reach level $i$ is at most $\lfloor k/2^{i-1}\rfloor$, leading the theorem. ∎
Therefore, each leader performs HEO-DFS at most $O(\log k)$ times. According to the analysis in Section 3, each HEO-DFS completes in $O(k\log\tau)$ time, which seems to imply that $\mathbf{GeneralDispersion}$ finishes in $O((\log k)\cdot(k\log\tau))=O(k\log^{2}k)$ time. However, this analysis does not take into account the length of the period during which leaders suspend their HEO-DFS. Thus, it is not clear whether a naive implementation of the strategy described above would achieve the dispersion in $O(k\log^{2}k)$ time. Following Shintaku et al.[14], we vary the speed of zombies chasing leaders based on a certain condition, which bounds the execution time by $O(k\log^{2}k)$ time.
We classify zombies based on two variables $\mathtt{level}_{L}$ and $\mathtt{level}_{S}$ that each zombie manages. For any zombie $z$, we call $z.\mathtt{level}_{L}$ and $z.\mathtt{level}_{S}$ the location level and swarm level of $z$. When a leader $z$ becomes a zombie, it initializes both $z.\mathtt{level}_{L}$ and $z.\mathtt{level}_{S}$ with its level, i.e.$z.\mathtt{level}$. Thereafter, a zombie $z$ copies the level of $\psi(\nu(z))$ to $z.\mathtt{level}_{L}$ and updates $z.\mathtt{level}_{S}$ to be $\max\{b.\mathtt{level}\mid b\in A_{Z}(\nu(z))\}$ in every $O(1)$ time steps. Since a zombie only chases a leader with an equal or greater level, $z.\mathtt{level}_{S}\leq z.\mathtt{level}_{L}$ always holds. We say that a zombie $z$ is strong if $z.\mathtt{level}_{S}=z.\mathtt{level}_{L}$; $z$ is weak otherwise.
We exploit the assumption that the agents are synchronous and let weak zombies move twice as frequently as strong zombies to chase a leader. As we will prove later, this difference in chasing speed results in a desirable property of $\mathbf{GeneralDispersion}$, namely that $\min(\{a.\mathtt{level}\mid a\in A_{AL}\}\cup\{z.\mathtt{level}_{L}\mid z\in A_{Z}\})$ is monotone non-decreasing and increases by at least one in every $O(k\log\tau)$ steps, where $A_{AL}$ is the set of active leaders and $A_{Z}$ is the set of zombies both in the whole graph, until $A_{AL}\cup A_{Z}$ becomes empty. Thus, by Lemma 3, $A_{AL}\cup A_{Z}$ becomes empty and the dispersion is achieved in $O(k(\log\tau)\cdot(\log k))=O(k\log^{2}k)$ steps.
@@
2       /*************** Slot 1 begins ***************/ Let $w=\nu(a)$. if $\exists b\in A_{L}(w)\cup A_{S}(w):a\prec b$  then
             $a.\mathtt{level}_{L}\leftarrow a.\mathtt{level}_{S}\leftarrow a.\mathtt{level}$ $a.\mathtt{leader}\leftarrow b.\mathtt{leader}$   // $a$ becomes a zombie and stops Algorithm 2
3
4      /*************** Slot 2 begins ***************/ if $A(w)\neq\{a\}$ then // $a$ is an active leader if $A(w)\neq\{a\}$
@@ 7                         Settle one zombie in $A_{Z}(w)$ at $w$
                  $\psi(w).\mathtt{parent}\leftarrow a.\mathtt{parent}$  // Initially, $a.\mathtt{parent}=\bot$
8
@@
10                   $(a.\mathtt{level},b.\mathtt{level})\leftarrow(a.\mathtt{level}+1,0)$ $\psi(w).\mathtt{parent}\leftarrow\bot$ $a.\mathtt{InitProbe}\leftarrow\mathbf{true}$
11            $(\psi(w).\mathtt{leader},\psi(w).\mathtt{level})\leftarrow(a.\mathtt{ID},a.\mathtt{level})$ if $a.\mathtt{InitProbe}=\mathbf{true}$ then // Initially, $a.\mathtt{InitProbe}=\mathbf{true}$
12                   $(\psi(w).\mathtt{next},\psi(w).\mathtt{checked},\psi(w).\mathtt{help},\psi(w).\mathtt{done})\leftarrow(\bot,-1,\bot,\mathbf{false})$ $a.\mathtt{InitProbe}\leftarrow\mathbf{false}$
            $\mathtt{Probe}(a)$   // See Algorithm 3
@@
14                   /*************** Slot 11 begins ***************/ if $\psi(w).\mathtt{next}=\bot$ then
                        $\psi(w).\mathtt{next}\leftarrow\psi(w).\mathtt{parent}$   // for backward move
15
16                  All agents in $A(w)\setminus\{\psi(w)\}$ move to $N(w,\psi(w).\mathtt{next})$ /*************** Slot 12 begins ***************/ $a.\mathtt{parent}\leftarrow a.\mathtt{pin}$ $a.\mathtt{InitProbe}\leftarrow\mathbf{true}$
17
18
@@
19 /*************** Slot 4 begins ***************/ Let $w=\nu(a)$. $\psi(w).\mathtt{next}(w)\leftarrow\begin{cases}\min P&\text{if\ }P\neq\emptyset\\ \bot&\text{otherwise},\end{cases}$
20    where $P=[0,\psi(w).\mathtt{checked}]\setminus\{b.\mathtt{pin}\mid b\in A_{S}(w)\setminus\{\psi(a)\}\}$ $b.\mathtt{help}\leftarrow\bot$ for all $b\in A_{S}(w)$ with $b\prec a$. Let all agents $b\in A_{S}(w)$ with $b\prec a$ go back to their homes if $\psi(w).\mathtt{next}\neq\bot\lor\psi(w).\mathtt{checked}=\delta_{w}-1$ then
21       /*************** Slot 5 begins ***************/ Execute $b.\mathtt{help}\leftarrow\bot$ for each $b\in A_{S}(w)\setminus\{\psi(w)\}$ Let all agents in $A_{S}(w)\setminus\{\psi(w)\}$ go back to their homes. $\psi(w).\mathtt{done}\leftarrow\mathbf{true}$
22else
23       /*************** Slot 6 begins ***************/ Let $\{a_{1},a_{2},\dots,a_{x}\}$ be the set of agents in $A(w)\setminus\{\psi(w)\}$ Let $\delta^{\prime}=\min(x,\delta_{w}-1-\psi(w).\mathtt{checked})$ Let $u_{i}=N(w,i+\psi(w).\mathtt{checked})$ for $i=1,2,\dots,\delta^{\prime}$ for each $a_{i}\in\{a_{1},a_{2},\dots,a_{\delta^{\prime}}\}$ in parallel do
24             $a_{i}$ moves to $u_{i}$. /*************** Slot 7 begins ***************/ if $(a_{i}.\mathtt{leader},a_{i}.\mathtt{level})=(\psi(u_{i}).\mathtt{leader},\psi(u_{i}).\mathtt{level})$ then
25                   $a_{i}.\mathtt{found}\leftarrow\mathbf{true}$ $\psi(u_{i}).\mathtt{help}\leftarrow a_{i}.\mathtt{pin}$
@@
29      /*************** Slot 8 begins ***************/ if $\exists i\in[1,\delta^{\prime}]:a_{i}.\mathtt{found}=\mathbf{false}$ then
@@
31      $\psi(w).\mathtt{checked}\leftarrow\psi(w).\mathtt{checked}+\delta^{\prime}$ Let all agents in $A_{S}(w)\setminus\{\psi(w)\}$ go back to their homes.
@@
/*************** Slot 3 begins ***************/ Move to $N(\nu(s),s.\mathtt{help})$ if $s.\mathtt{help}\neq\bot$.
@@
32 /*************** Slot 9 begins ***************/ $(z.\mathtt{level}_{L},z.\mathtt{level}_{S})\leftarrow(\psi(w).\mathtt{level},\max\{z^{\prime}.\mathtt{level}\mid z^{\prime}\in A_{Z}(\nu(z)\})$ if $A_{L}(\nu(z))=\emptyset$ and $z$ is a weak zombie then
@@
34/*************** Slot 10 begins ***************/ if $A_{L}(\nu(z))=\emptyset$ then
@@
4.2 Details
In $\mathbf{GeneralDispersion}$, we group every 12 time steps into one unit, with each unit consisting of twelve slots. In other words, time steps $0,1,2,\dots$ are classified into twelve slots. Specifically, each time step $t\geq 0$ is assigned to slot $(t\bmod 12)+1$. For example, time step 26 is in slot 3, and time step 47 is in slot 12. Dividing all time steps into twelve slots helps to reduce the interference of multiple HEO-DFSs and allows us to set different “chasing speeds” for weak and strong zombies. Table 2 summarizes the roles of each slot. Note that each agent needs to manage an $O(1)$-bit variable to identify the slot of the current time step, but for simplicity, the process related to its update is not included in the pseudocode because it can be implemented in a naive way.
The pseudocode for the $\mathbf{GeneralDispersion}$ algorithm is shown in Algorithms 2, 3, 4, and 5. In slots 1, 2, 4–8, 11, and 12, agents operate only under the instruction of a leader. Algorithms 2 and 3 define how each leader $a$ operates and gives instructions in those slots. Algorithm 4 defines the behavior of settlers in slot 3. Algorithm 5 specifies the behavior of zombies in slots 9 and 10.

### v3

@@ 4.1 Overview
A level-up event requires one leader $a$ and one zombie $b$ with the same level. That zombie $b$ will get level 0. Thereafter, $b$ never triggers a level-up event again because the level of a leader is monotonically non-decreasing starting from level 1. Therefore, for any $i\geq 1$, the number of agents that can reach level $i$ is at most $\lfloor k/2^{i-1}\rfloor$, leading the lemma. ∎
Therefore, each leader performs HEO-DFS at most $O(\log k)$ times. According to the analysis in Section 3, each HEO-DFS completes in $O(k\log\tau)$ time, which seems to imply that $\mathbf{GeneralDisp}$ finishes in $O((\log k)\cdot(k\log\tau))=O(k\log^{2}k)$ time. However, this analysis does not take into account the length of the period during which leaders suspend their HEO-DFS. Thus, it is not clear whether a naive implementation of the strategy described above would achieve the dispersion in $O(k\log^{2}k)$ time. Following Shintaku et al. [16], we vary the speed of zombies chasing leaders based on a certain condition, which bounds the execution time by $O(k\log^{2}k)$ time.
We give zombies different chasing speeds as follows. First, we classify zombies based on two variables $\mathtt{level}_{L}$ and $\mathtt{level}_{S}$ that each zombie manages. For any zombie $z$, we call $z.\mathtt{level}_{L}$ and $z.\mathtt{level}_{S}$ the location level and swarm level of $z$. When a leader $z$ becomes a zombie, it initializes both $z.\mathtt{level}_{L}$ and $z.\mathtt{level}_{S}$ with its level, i.e.$z.\mathtt{level}$. Thereafter, a zombie $z$ copies the level of $\psi(\nu(z))$ to $z.\mathtt{level}_{L}$ and updates $z.\mathtt{level}_{S}$ to be $\max\{b.\mathtt{level}\mid b\in A_{Z}(\nu(z))\}$ in every $O(1)$ time steps. Since a zombie only chases a leader with an equal or greater level, $z.\mathtt{level}_{S}\leq z.\mathtt{level}_{L}$ always holds. We say that a zombie $z$ is strong if $z.\mathtt{level}_{S}=z.\mathtt{level}_{L}$; $z$ is weak otherwise. Then, we exploit the assumption that the agents are synchronous and let weak zombies move twice as frequently as strong zombies to chase a leader. As we will prove later, this difference in chasing speed results in a desirable property of $\mathbf{GeneralDisp}$, namely that $\min(\{a.\mathtt{level}\mid a\in A_{AL}\}\cup\{z.\mathtt{level}_{L}\mid z\in A_{Z}\})$ is monotone non-decreasing and increases by at least one in every $O(k\log\tau)$ steps, where $A_{AL}$ is the set of active leaders and $A_{Z}$ is the set of zombies both in the whole graph, until $A_{AL}\cup A_{Z}$ becomes empty. Thus, by Lemma 3, $A_{AL}\cup A_{Z}$ becomes empty and the dispersion is achieved in $O(k\log\tau\cdot\log k)=O(k\log^{2}k)$ steps.
@@
Figure 2: The behavior of explorers when their leader invokes $\mathtt{Probe}()$ at the center node $w$ in $\mathbf{GeneralDisp}$. A black circle, triangle, and rectangle represent a leader, a zombie, and a settler, respectively. The integers in the top left figure represents port numbers.
4.2 Time Slots
In $\mathbf{GeneralDisp}$, we group every 12 time steps into one unit, with each unit consisting of twelve slots. In other words, time steps $0,1,2,\dots$ are classified into twelve slots. Specifically, each time step $t\geq 0$ is assigned to slot $(t\bmod 12)+1$. For example, time step 26 is in slot 3, and time step 47 is in slot 12. Dividing all time steps into twelve slots helps to reduce the interference of multiple HEO-DFSs and allows us to set different “chasing speeds” for weak and strong zombies. Table 2 summarizes the roles of each slot.
Essentially, slots 1–2 are designated for leader election (i.e., group merging), slots 3–8 for probing, slots 9–10 for zombie chasing, and slots 11–12 for forward and backward movement in DFS traversal. It is important to note that settlers always stay at their home during slots 1–2 and 9–12. Hence, once a settler leaves its home, it returns within $O(1)$ steps. This is not the case in $\mathbf{RootedDisp}$, where a settler in helping mode does not return home until its leader completes the probing process. In $\mathbf{GeneralDisp}$, this frequent return home enables leaders to detect collisions with other groups: if a leader enters another group’s territory, it will certainly notice the intrusion during the next slot 1, as it encounters a settler from that group.
Thus, the probing process in $\mathbf{GeneralDisp}$ slightly differs from that in $\mathbf{RootedDisp}$. Consider a leader $a_{l}$ starting the probing process at a node $w$ (refer to Figure 2). The objective here is to identify any neighboring node of $w$ that lies outside $a_{l}$’s territory, if such exists. During slots 6–7, explorers at $w$ visit its neighbors and return in parallel. If an explorer $b$ encounters a settler $s$ at a node $u\in N(w)$ in slot 6, $b$ does not bring $s$ back to $w$ in slot 7. Instead, $b$ requests $s$ to enter helping mode, wherein $s$ records the port number to $w$ in the variable $s.\mathtt{help}\in\mathbb{N}\cup{\bot}$ (A settler $s$ is in helping mode if and only if $s.\mathtt{mode}\neq\bot$). The helping settler $s$ moves to $w$ via port $s.\mathtt{help}$ in the subsequent slot 3, joins the probing in slots 6-7, and returns to its home $u$ again in slot 8. Leader $a_{l}$ expects that the exact $\psi(w).\mathtt{checked}$ helping settlers arrives at $w$ in each slot 3. If this does not occur, $a_{l}$ detects a non-territorial neighbor in slot 4 by checking the $\mathtt{pin}$ variable of the helping settlers at $w$. Similar to $\mathbf{RootedDisp}$’s probing process, the total number of explorers and helping settlers at $w$ doubles until such a neighbor is detected, concluding the process in $O(\log\tau)$ steps. Subsequently, $a_{l}$ reverts the helping settlers at $w$ to non-helping mode by setting their $\mathtt{help}$ variable to $\bot$ and instructs them to return to their homes in slot 5.
As mentioned earlier, we differentiate the chasing speed of weak zombies and strong zombies. Specifically, weak zombies move in both slots 9 and 10, while strong zombies move only in slot 10.
Before presenting the detailed implementation of $\mathbf{GeneralDisp}$ including pseudocodes in Section 4.3, we give a proof sketch here for the following main theorem.
Theorem 2.
In the general setting, there exists an algorithm that solves the dispersion problem within $O(k\log\tau\cdot\log k)$ time using $O(\log(k+\Delta))$ bits of space per agent.
Proof Sketch.
It suffices to show that $A_{AL}\cup A_{Z}$ becomes empty within $O(k\log\tau\cdot\log k)$ steps, at which point every agent is either a waiting leader or a settler, thereby achieving dispersion. We obtain this bound from Lemma 3 and the fact that $\alpha=\min(\{a.\mathtt{level}\mid a\in A_{AL}\}\cup\{z.\mathtt{level}_{L}\mid z\in A_{Z}\})$ increases by at least one in every $O(k\log\tau)$ time steps unless $A_{AL}\cup A_{Z}$ becomes empty (Lemma 6 in Section 4.4). We can prove Lemma 6 roughly as follows. Suppose $\alpha=i$. First, all weak zombies at location level $i$ vanish within $O(k)$ steps as they move faster than leaders and strong zombies, eventually encountering a leader, higher-level settlers or stronger zombies. Thereafter, no new weak zombies at location level $i$ are created. Subsequently, without weak zombies at location level $i$, waiting leaders at level $i$ do not resume active HEO-DFS without increasing its level, which leads to the disappearance of active leaders at level $i$ within $O(k\log\tau)$ steps. Finally, strong zombies chasing leaders at level $i$ catch up to those leaders within $O(k)$ steps or find higher-level settlers, resulting in the increase of their location level. Hence, all zombies with location level $i$ and active leaders at level $i$ are eliminated within $O(k\log\tau)$ steps. ∎
@@ 4.2 Time Slots
2       /*************** Slot 1 begins ***************/
3       Let $w=\nu(a)$.
4       if $\exists b\in A_{L}(w)\cup A_{S}(w):a\prec b$  then
5             $a.\mathtt{level}_{L}\leftarrow a.\mathtt{level}_{S}\leftarrow a.\mathtt{level}$
             $a.\mathtt{leader}\leftarrow b.\mathtt{leader}$
              // $a$ becomes a zombie and stops Algorithm 2
6
7      /*************** Slot 2 begins ***************/
8       if $A(w)\neq\{a\}$ then // $a$ is an active leader if $A(w)\neq\{a\}$
@@ 11                         Settle one zombie in $A_{Z}(w)$ at $w$
12
                  $\psi(w).\mathtt{parent}\leftarrow a.\mathtt{parent}$
                    // Initially, $a.\mathtt{parent}=\bot$
13
@@
15                   $(a.\mathtt{level},b.\mathtt{level})\leftarrow(a.\mathtt{level}+1,0)$
16                   $\psi(w).\mathtt{parent}\leftarrow\bot$
17                   $a.\mathtt{InitProbe}\leftarrow\mathbf{true}$
18
19            $(\psi(w).\mathtt{leader},\psi(w).\mathtt{level})\leftarrow(a.\mathtt{ID},a.\mathtt{level})$
20             if $a.\mathtt{InitProbe}=\mathbf{true}$ then // Initially, $a.\mathtt{InitProbe}=\mathbf{true}$
21                   $(\psi(w).\mathtt{next},\psi(w).\mathtt{checked},\psi(w).\mathtt{help},\psi(w).\mathtt{done})\leftarrow(\bot,-1,\bot,\mathbf{false})$
22                   $a.\mathtt{InitProbe}\leftarrow\mathbf{false}$
            $\mathtt{Probe}(a)$
              // See Algorithm 3
@@
24                   /*************** Slot 11 begins ***************/
25                   if $\psi(w).\mathtt{next}=\bot$ then
                        $\psi(w).\mathtt{next}\leftarrow\psi(w).\mathtt{parent}$
                          // for backward move
26
27                  All agents in $A(w)\setminus\{\psi(w)\}$ move to $N(w,\psi(w).\mathtt{next})$
28                   /*************** Slot 12 begins ***************/
29                   $a.\mathtt{parent}\leftarrow a.\mathtt{pin}$
30                   $a.\mathtt{InitProbe}\leftarrow\mathbf{true}$
31
32
33
@@
34 /*************** Slot 4 begins ***************/
35 Let $w=\nu(a)$.
36 $\psi(w).\mathtt{next}(w)\leftarrow\begin{cases}\min P&\text{if\ }P\neq\emptyset\\ \bot&\text{otherwise},\end{cases}$
37    where $P=[0,\psi(w).\mathtt{checked}]\setminus\{b.\mathtt{pin}\mid b\in A_{S}(w)\setminus\{\psi(a)\}\}$
38 $b.\mathtt{help}\leftarrow\bot$ for all $b\in A_{S}(w)$ with $b\prec a$.
39 Let all agents $b\in A_{S}(w)$ with $b\prec a$ go back to their homes
40 if $\psi(w).\mathtt{next}\neq\bot\lor\psi(w).\mathtt{checked}=\delta_{w}-1$ then
41       /*************** Slot 5 begins ***************/
42       Execute $b.\mathtt{help}\leftarrow\bot$ for each $b\in A_{S}(w)\setminus\{\psi(w)\}$
43       Let all agents in $A_{S}(w)\setminus\{\psi(w)\}$ go back to their homes.
44       $\psi(w).\mathtt{done}\leftarrow\mathbf{true}$
45
46else
47       /*************** Slot 6 begins ***************/
48       Let $\{a_{1},a_{2},\dots,a_{x}\}$ be the set of agents in $A(w)\setminus\{\psi(w)\}$
49       Let $\Delta^{\prime}=\min(x,\delta_{w}-1-\psi(w).\mathtt{checked})$
50       Let $u_{i}=N(w,i+\psi(w).\mathtt{checked})$ for $i=1,2,\dots,\Delta^{\prime}$
51       for each $a_{i}\in\{a_{1},a_{2},\dots,a_{\Delta^{\prime}}\}$ in parallel do
52             $a_{i}$ moves to $u_{i}$.
53             /*************** Slot 7 begins ***************/
54             if $(a_{i}.\mathtt{leader},a_{i}.\mathtt{level})=(\psi(u_{i}).\mathtt{leader},\psi(u_{i}).\mathtt{level})$ then
55                   $a_{i}.\mathtt{found}\leftarrow\mathbf{true}$
56                   $\psi(u_{i}).\mathtt{help}\leftarrow a_{i}.\mathtt{pin}$
@@ 48       Let $\{a_{1},a_{2},\dots,a_{x}\}$ be the set of agents in $A(w)\setminus\{\psi(w)\}$
60
61      /*************** Slot 8 begins ***************/
62       if $\exists i\in[1,\Delta^{\prime}]:a_{i}.\mathtt{found}=\mathbf{false}$ then
@@
65       Let all agents in $A_{S}(w)\setminus\{\psi(w)\}$ go back to their homes.
66
@@
67 /*************** Slot 3 begins ***************/
Move to $N(\nu(s),s.\mathtt{help})$ if $s.\mathtt{help}\neq\bot$.
@@
68 /*************** Slot 9 begins ***************/
69 $(z.\mathtt{level}_{L},z.\mathtt{level}_{S})\leftarrow(\psi(w).\mathtt{level},\max\{z^{\prime}.\mathtt{level}\mid z^{\prime}\in A_{Z}(\nu(z)\})$
70 if $A_{L}(\nu(z))=\emptyset$ and $z$ is a weak zombie then
@@
72/*************** Slot 10 begins ***************/
73 if $A_{L}(\nu(z))=\emptyset$ then
@@
4.3 Detail Implementation of General Dispersion
The pseudocode for the $\mathbf{GeneralDisp}$ algorithm is shown in Algorithms 2, 3, 4, and 5. In slots 1, 2, 4–8, 11, and 12, agents operate only under the instruction of a leader. Algorithms 2 and 3 define how each leader $a$ operates and gives instructions in those slots. Algorithm 4 defines the behavior of settlers in slot 3. Algorithm 5 specifies the behavior of zombies in slots 9 and 10. Note that each agent needs to manage an $O(1)$-bit variable to identify the slot of the current time step, but for simplicity, the process related to its update is not included in the pseudocode because it can be implemented in a naive way.